# Word Order in English-Japanese Simultaneous Interpretation: Analyses and Evaluation using Chunk-wise Monotonic Translation

### v1

@@ 4.2.2 Factors that Make CMT Sentences Longer
In addition to the above characteristics, there were many repetitions observed in particularly long sentences. To verify this, we further analyzed particularly long and short sentences, chosen based on the length ratio of CMT sentences to offline ones. The long and short sentences were defined as those with a length ratio greater/smaller than the average 0.5 standard deviations (avg.=1.39, SD=0.43). We subjectively judged if those sentences contained many repetitions. We also identified the sentences whose offline translations were short.
Table 5 shows that long sentences contained more repetitions than short ones. The offline translation sentences were short, probably because they were originally subtitles, for which limited space was allowed. We also checked it quantitatively using the number of repeat tags assigned to the sentences, and found that the frequency of repetition tags was higher in long sentences (Table 5).
@@ 4.2.3 Omission in SI Sentences
To find techniques to shorten translations, we analyzed SI sentences in NAIST-SIC. Based on the length ratio of SI sentences to offline ones, we defined SI sentences that potentially involved reasonable omissions (omission; the ratio ), and SI sentences that may not fully convey the meaning of the source sentences (undertranslation; the ratio ), following the criteria in Higashiyama et al. (2023).
We expected to find some trends (e.g., part-of-speech) in phrases that were omitted, but we was not able to. In addition, we found a certain number of unacceptable translations in both categories (43.12% and 60.00% for omission and undertranslation, respectively). The results suggest that human simultaneous interpreters judge the importance of the phrases based on context and decide whether to translate them, which are sometimes correct and sometimes not.
@@ 4.2.4 Factors that Make Monotonic Translation Difficult
With the help of tags annotated to source and CMT sentences, we analyzed factors that make monotonic translation difficult. Table 6 shows the number of source phrases that were repeated, deferred, and omitted. The values are based on the number of rel-repeat, defer_d#, and omit tags, respectively. We counted the number of relation tags for repeat and defer because the span tags for the two operations were assigned to both source and CMT sentences. The results show that the translators used repeat most frequently, followed by defer and omit, as they were instructed (see Section 3.1).
@@
For these phrases, we explored what makes monotonic translations difficult. Our analyses revealed that most cases of repeat and defer occurred because of particular grammatical structures. Table 7 lists major structures along with their frequencies in the data and examples. In these structures, a phrase in a chunk is typically a dependent of a phrase in the preceding chunk. In the example of a post-modifier (Table 7), the relative pronoun clause is a dependent of the noun phrase a device, which is in a preceding chunk. When phrases with a dependency relation exist across multiple chunks, CMT would be difficult because Japanese is a strongly head-final language. The examples in Table 7 show how human translators address these structures by repeating or deferring some phrases in the subsequent chunks.
Prepositions, post-modifiers, and dependent clauses have also been identified as syntactic factors that affect interpreters’ word order decisions in Cai et al. (2020). Human interpreters find those structures challenging for SI and adopt a strategy to maintain the word order of the source language.
@@
In addition, we observed that inappropriate segmentation was addressed through repeating and deferring the phrases. Most of the inappropriate segmentation was found in phrasal verbs, -ing forms of verbs, and to-infinitives.
In SI data, we also found human interpreters repeat phrases to maintain the word order of the source language. For example, in the example in Table 1, a noun modified by a preposition phrase is repeated.
@@
In addition, Okamura and Yamada (2023) reported that the order of the chunks is swapped about once on average in an SI sentence. These things suggest that human interpreters address the word order differences that make monotonic translation difficult through repeating and deferring some phrases.
5 Evaluation Using CMT Data
To investigate the impact of using CMT sentences for evaluating the translation quality, we evaluated the output from existing ST and simulST models using NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as existing test sets.
@@ 5.1 Data
We used the following datasets as the reference for the automatic evaluation metrics:
n-cmt: CMT sentences from NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset
si_hum: SI sentences from NAIST-SIC, which were manually aligned to the source speech
si_auto: SI sentences from NAIST-SIC-Aligned, which were aligned automatically
@@ 5.2 Speech Translation Models
We used three existing ST and simulST models:
@@
simulST_si_offline: a simulST model trained on both offline and SI data Ko et al. (2023)
All the models were built by connecting two pre-trained models, HuBERT-Large Hsu et al. (2021) for their speech encoder and the decoder of mBART50 Tang et al. (2020) for their text decoder. The encoder and decoder were connected via Inter-connection Nishikawa and Nakamura (2023) and a length adapter Tsiamas et al. (2022). Both SimulST models used bilingual prefix pairs extracted using Bilingual Prefix Alignment Kano et al. (2022) for the model training and employed a decoding policy called local agreement Liu et al. (2020). For ST_offline, we used the model with checkpoint averaging (Inter-connection + Ckpt Ave. in Fukuda et al. (2023)). For simulST_offline and simulST_si_offline, we used the models which satisfy the task requirement of the simultaneous track in IWSLT 2023 Evaluation Campaign999https://iwslt.org/2023/simultaneous, latency measured by Average Lagging Ma et al. (2019) seconds (Offline FT and Mixed FT + Style in Ko et al. (2023), respectively).
5.3 Metrics
We evaluated the translation quality of the output from the ST and simulST models (see Section 5.2) using BLEU101010BLEU was calculated using sacreBLEU. Post (2018) https://github.com/mjpost/sacrebleu, BLEURT111111https://github.com/google-research/bleurt Sellam et al. (2020), COMET121212https://github.com/Unbabel/COMET Rei et al. (2020), and BERTScore131313https://github.com/Tiiiger/bert_score Zhang et al. (2020). BERTScore was calculated using bert-base-multilingual-cased. We used the four datasets described in Section 5.1 as the reference.
5.4 Evaluation Results
@@
| ST_offline | 14.487 | 8.856 | 8.637 | 17.775 | 0.553 | 0.447 | 0.414 | 0.538 | 0.838 | 0.797 | 0.781 | 0.833 |
| simulST_offline | 15.406 | 8.446 | 7.773 | 17.907 | 0.556 | 0.442 | 0.406 | 0.531 | 0.826 | 0.780 | 0.763 | 0.821 |
| simulST_si_offline | 15.982 | 12.031 | 11.020 | 13.191 | 0.567 | 0.493 | 0.460 | 0.519 | 0.807 | 0.774 | 0.761 | 0.789 |
@@
| simulST_si_offline | 0.817 | 0.764 | 0.746 | 0.759 | 0.784 | 0.766 | 0.760 | 0.757 | 0.800 | 0.764 | 0.752 | 0.757 |
@@
Table 8 shows the results of quality evaluation metrics across ST and simulST models. When the translation quality was evaluated using BLEU with n-cmt as the reference, simulST_si_offline achieved the highest score. On the SI-based test sets (i.e., si_hum and si_auto), simulST_si_offline also had the highest score. On the offline-based test set, in contrast, models trained on only offline data achieved much higher score than simulST_si_offline. The same tendencies were observed in BLEURT and BERTScore. These results suggest that the model trained on both SI and offline data generated more SI-like translations, and that such models should be evaluated using a reference closer to SI sentences. In addition, using an offline-based test set might underestimate the performance of the model trained on both SI and offline data.
@@
Comparing n-cmt, si_hum, and si_auto, the scores were highest for n-cmt, followed by si_hum and si_auto on all metrics and models. Because si_hum is based on SI sentences generated by human simultaneous interpreters, some content in a source speech might be omitted or inadequately translated (under-translation). SI sentences in si_auto, which were automatically created based on human SI sentences, might contain less source speech content than those in si_hum due to the alignment and filtering techniques applied (see Zhao et al., 2024). In fact, BERTScore precision was higher than recall on n-cmt, in which there are almost no omissions, while the recall was higher than the precision on si_auto and the precision and the recall were almost equal on si_hum. These results suggest that the existing SI-based test sets Ko et al. (2023); Zhao et al. (2024) underestimate the model performance.
@@
However, the results on COMET were different from those on other metrics (Table 8). On all four test sets, ST_offline achieved the highest score, followed by simulST_offline and simulST_si_offline. One of the possible reasons is that COMET uses source sentences to calculate scores.
@@
To examine the impact of source sentences, we also calculated the reference-free COMET-QE using wmt22-cometkiwi-da and got similar results (0.813, 0.798, 0.766 for ST_offline, simulST_offline, simulST_si_offline, respectively). We further calculated COMET-QE for n-cmt and offline, regarding them as oracle data, and found that n-cmt had a higher score than offline (n-cmt: 0.832, offline: 0.812). Because some translation sentences in offline include under-translation, these results suggest that the COMET score tends to become higher when more content in the source sentences is covered in the target sentences. This feature does not fit the nature of SI, where human interpreters use sophisticated strategies (see He et al., 2016; Cai et al., 2020, for example). We need to carefully interpret the COMET score when we use it for evaluating simulST models.
@@ 6 Conclusion
This paper focuses on monotonic translations in English-Japanese SI. Our analyses revealed some syntactic factors that make monotonic translations difficult and that human interpreters/translators address these challenges by repeating or deferring some phrases in the source language in the subsequent chunks. The grammatical structures that might cause delay would be useful information for developing segmentation or decoding policies of simultaneous machine translation systems.
We also evaluated the output from the existing ST and simulST models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing SI-based and offline-based test sets. The results suggest that monotonic translation data would better evaluate simulST models trained using SI data, and that the existing test sets underestimate the performance of such models. A future direction would be using monotonic translation data for developing simulST models.

### v2

@@ 4.2.2 Factors that Lengthen CMT Sentences
In addition to the above characteristics, we observed many repetitions in particularly long sentences. To verify this, we further analyzed particularly long and short sentences, chosen based on the length ratio of the CMT sentences to the offline ones. The long and short sentences were defined as those with a length ratio greater/smaller than the average 0.5 standard deviations (avg.=1.39, SD=0.43). We subjectively judged whether these sentences contained many repetitions. We also identified sentences whose offline translations were short.
Table 5 shows that long sentences contained more repetitions than short ones. The offline translation sentences were short, probably because they were originally subtitles, for which limited space was allowed. We also quantitatively checked them using the number of assigned repeat tags and found that the frequency of repetition tags was higher in long sentences (Table 5).
@@ 4.2.3 Omission in SI Sentences
To find techniques for shortening translations, we analyzed the SI sentences in NAIST-SIC. Based on the length ratio of SI sentences to the offline ones, we defined SI sentences that might have reasonable omissions (omission; ratio ) and SI sentences that probably failed to fully convey the meaning of the source sentences (undertranslation; ratio ), following the criteria in Higashiyama et al. (2023).
Although we expected to identify some trends (e.g., part-of-speech) in the phrases that were omitted, we did not do so. In addition, we found a certain number of unacceptable translations in both categories (43.12% and 60.00% for omission and undertranslation, respectively). The results suggest that human simultaneous interpreters judge the importance of phrases based on context and decide whether to translate them; some judgements are correct, and some are not.
@@ 4.2.4 Factors that Make Monotonic Translation Difficult
With the help of tags annotated to the source and CMT sentences, we analyzed the factors that make monotonic translation difficult. Table 6 shows the number of source phrases that were repeated, deferred, or omitted. The values are based on the number of rel-repeat, defer_d#, and omit tags. We counted the relation tags for repeat and defer because the span tags for those two operations were assigned to both the source and CMT sentences. The results show that the translators used repeat most frequently, followed by defer and omit, as they were instructed (see Section 3.1).
@@
For these phrases, we explored what makes monotonic translations difficult. Our analyses revealed that most cases of repeat and defer were caused by particular grammatical structures. Table 7 lists the major structures along with their frequencies in the data and examples. In these structures, a phrase in a chunk is typically a dependent of a phrase in the preceding chunk. In the example of a post-modifier (Table 7), the relative pronoun clause is a dependent of the noun phrase a device, which is in a preceding chunk. When phrases with a dependency relation exist across multiple chunks, CMT is difficult because Japanese is a strongly head-final language. The examples in Table 7 show how human translators address these structures by repeating or deferring some phrases in subsequent chunks.
Prepositions, post-modifiers, and dependent clauses have also been identified as syntactic factors that affect interpreters’ word order decisions in Cai et al. (2020). Human interpreters find these structures challenging for SI and adopt a strategy to maintain the word order of the source language.
@@
In addition, we observed that inappropriate segmentation was addressed by repeating and deferring the phrases. Most inappropriate segmentation was found in phrasal verbs, verbal gerunds, and to-infinitives.
In the SI data, we also found that human interpreters repeat phrases to maintain the word order of the source language. For example, in the example in Table 1, a noun modified by a preposition phrase is repeated:
@@
In addition, Okamura and Yamada (2023) reported that the order of the chunks is shuffled about once on average in an SI sentence. These things suggest that human interpreters address the word order differences that make monotonic translation difficult by repeating and deferring some phrases.
5 Evaluation Using CMT sentences
To investigate the impact of using CMT sentences for evaluating translation quality, we evaluated the output from existing ST and simulST models using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as existing test sets.
@@ 5.1 Data
We used the following four datasets as references for the automatic evaluation metrics:
n-cmt: CMT sentences from the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset
si_hum: SI sentences from NAIST-SIC, manually aligned to the source speech
si_auto: SI sentences from NAIST-SIC-Aligned, aligned automatically
@@
Because si_auto was created by applying automatic alignment and filtering techniques to SI sentences in si_hum, it may contain alignment errors. In addition, SI sentences in si_auto tend to be shorter than those in si_hum (see Table 4). Therefore, we used the two SI-based datasets for our evaluation.
@@ 5.2 Speech Translation Models
We used three existing models (i.e., one ST and two simulST models):
@@
simulST_si_offline: a simulST model trained on both offline and SI data Ko et al. (2023).
All the models were built by connecting two pre-trained models, HuBERT-Large Hsu et al. (2021) for their speech encoder and the decoder of mBART50 Tang et al. (2020) for their text decoder. The encoder and decoder were connected by Inter-connection Nishikawa and Nakamura (2023) and a length adapter Tsiamas et al. (2022). Both SimulST models used bilingual prefix pairs extracted using Bilingual Prefix Alignment Kano et al. (2022) for the model training and employed a decoding policy called local agreement Liu et al. (2020). For ST_offline, we used a model with checkpoint averaging (Inter-connection + Ckpt Ave. in Fukuda et al. (2023)). For simulST_offline and simulST_si_offline, we used the models that satisfy the task requirement of the simultaneous track in the IWSLT 2023 Evaluation Campaign111111https://iwslt.org/2023/simultaneous, latency measured by Average Lagging Ma et al. (2019) seconds (Offline FT and Mixed FT + Style in Ko et al. (2023), respectively).
@@
| ST_offline | 14.487 | 8.856 | 8.637 | 17.775 | 0.553 | 0.447 | 0.414 | 0.538 | 0.838 | 0.797 | 0.781∗1 | 0.833 |
| simulST_offline | 15.406† | 8.446† | 7.773† | 17.907 | 0.556 | 0.442 | 0.406 | 0.531 | 0.826 | 0.780 | 0.763 | 0.821 |
| simulST_si_offline | 15.982† | 12.031† | 11.020† | 13.191† | 0.567 | 0.493∗1 | 0.460∗1 | 0.519 | 0.807∗2 | 0.774∗3 | 0.761 | 0.789∗2 |
@@
| simulST_si_offline | 0.817∗1 | 0.764∗1 | 0.746∗1 | 0.759∗2 | 0.784∗1 | 0.766∗1 | 0.760∗1 | 0.757∗2 | 0.800∗1 | 0.764∗1 | 0.752∗1 | 0.757∗2 |
5.3 Metrics
We evaluated the translation quality of the output from the ST and simulST models (see Section 5.2) using BLEU121212BLEU was calculated using sacreBLEU. Post (2018) https://github.com/mjpost/sacrebleu, BLEURT131313https://github.com/google-research/bleurt Sellam et al. (2020), COMET141414https://github.com/Unbabel/COMET Rei et al. (2020), and BERTScore151515https://github.com/Tiiiger/bert_score Zhang et al. (2020). BERTScore was calculated using bert-base-multilingual-cased. We used the four datasets described in Section 5.1 as references.
5.4 Evaluation Results
Table 8 shows the results of the quality evaluation metrics across the ST and simulST models. For the BLEU scores, we conducted paired significance tests using paired bootstrap resampling Koehn (2004). We specified ST_offline as the baseline for the significance tests. For the other scores, we conducted a one-way ANOVA, followed by Tukey’s multiple comparisons test.
@@ 5.4 Evaluation Results
When the translation quality was evaluated using BLEU with n-cmt as the reference, simulST_si_offline achieved the highest score. On the SI-based test sets (i.e., si_hum and si_auto), simulST_si_offline also had the highest score. On the offline-based test set, in contrast, the models trained on only offline data achieved much higher scores than simulST_si_offline. The same tendencies were observed in BLEURT and BERTScore. These results suggest that the models trained on both SI and offline data generated more SI-like translations, and such models perhaps should be evaluated using a reference closer to SI sentences. In addition, using an offline-based test set might underestimate the performance of models trained on both SI and offline data.
@@
Comparing n-cmt, si_hum, and si_auto, the scores were highest for n-cmt, followed by si_hum and si_auto on all the metrics and models. Because si_hum is based on SI sentences generated by human simultaneous interpreters, some content in the source speech might be omitted or inadequately translated (under-translation). SI sentences in si_auto, which were automatically created based on human SI sentences, might contain less source speech content than those in si_hum due to the alignment and filtering techniques applied (see Zhao et al., 2024). In fact, BERTScore precision was higher than recall on n-cmt, in which there were almost no omissions, while recall was higher than precision on si_auto and precision and recall were almost equal on si_hum. These results indicate the possibility that the existing SI-based test sets Ko et al. (2023); Zhao et al. (2024) underestimate the model performance.
@@
However, the COMET results were different from those on the other metrics (Table 8). On all four test sets, ST_offline achieved the highest score, followed by simulST_offline and simulST_si_offline. One possible reason is that COMET uses source sentences to calculate its scores.
@@
To examine the impact of the source sentences, we also calculated a reference-free COMET-QE using wmt22-cometkiwi-da and got similar results (0.813, 0.798, and 0.766 for ST_offline, simulST_offline, and simulST_si_offline, respectively). We further calculated COMET-QE for n-cmt and offline, regarding them as oracle data, and found that n-cmt had a higher score than offline (n-cmt: 0.832, offline: 0.812). Because some translation sentences in offline are under-translated, these results suggest that the COMET scores tend to become high when more content in the source sentences is covered in the target sentences. This feature does not fit the nature of SI, where human interpreters use sophisticated strategies (see He et al., 2016; Cai et al., 2020, for example). We need to carefully interpret COMET scores when we use them for evaluating simulST models.
@@ 6 Conclusion
This paper focused on monotonic translations in English-Japanese SI. Our analyses revealed some grammatical structures that make monotonic translations difficult and that human interpreters/translators address these challenges by repeating or deferring some phrases in source language in the subsequent chunks. The grammatical structures that might cause delays would be useful information for developing segmentation or decoding policies for simultaneous machine translation systems. One possible direction would be predicting whether a phrase in a chunk is the head of a phrase in subsequent chunks.
We also evaluated the output from the existing ST and simulST models on the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing SI-based and offline-based test sets. The BLEU, BLEURT, and BERTScore results supported using CMT sentences for evaluating simulST models trained using SI data, although the results with COMET were different. Further analysis across various evaluation metrics is necessary. Analyzing how the source and target sentences are aligned monotonically on different types of translations (e.g., Han et al., 2021) would also be useful.
This paper investigated the impact of using CMT sentences for evaluation purposes. A future study would involve using monotonic translation sentences for developing simulST models Sakai et al. (2024)161616Published around the same time as the submission of this paper.. It could potentially address the problem that simulST models trained using SI sentences suffered from under-translation Ko et al. (2023). However, CMT sentences tend to be long. Investigating the trade-offs between longer CMT sentences and the potential cognitive load on listeners/readers might provide further insights.